Determine the molar solubility of BaF2 in a solution containing 0.0750 M LiF. Ksp (BaF2) = 1.7
Answer #1
at equilibrium :
[Ba2+]= x
[F-]= 2x + 0.0750

Ksp = [Ba2+][F-]^2 = (x) ( 2x + 0.0750)^2

x = molar solubility =3.0 x 10^-4 M
Answer #2
example
Determine the molar solubility of BaF2 in a solution containing 0.0750 M LiF. Ksp (BaF2) = 1.7 × 10-6.?

ans
at equilibrium :
[Ba2+]= x
[F-]= 2x + 0.0750

Ksp = [Ba2+][F-]^2 = (x) ( 2x + 0.0750)^2

x = molar solubility =3.0 x 10^-4 M
Answer #3
at equilibrium :
[Ba2+]= x
[F-]= 2x + 0.0750

Ksp = [Ba2+][F-]^2 = (x) ( 2x + 0.0750)^2

x = molar solubility =3.0 x 10^-4 M